The length of a right triangle is 14 centimeters shorter than the other leg, and the hypotenuse is 2 centimeters longer then the longer leg. Find the lengths of all three sides.
Answer by TimothyLamb(4379)

---
x = long leg
y = short leg
---
y = x - 14
h = x + 2
---
h = sqrt( xx + yy )
hh = xx + yy
(x + 2)(x + 2) = xx + (x - 14)(x - 14)
xx + 4x + 4 = xx + xx - 28x + 196
xx - 32x + 192 = 0
---
the above quadratic equation is in standard form, with a=1, b=-32, and c=192
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 -32 192
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 24
x = 8
---
h = sqrt( 24*24 + 8*8 )
h = 25.298221281347036
---
answer:
x = long leg = 24 cm
y = short leg = 8 cm
h = hypotenuse = 25.2 cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php